if the measure of a regular polygons interior angles is 156 degrees, how many sides does the polygon have?
Answer by Alan3354(69443)

if the measure of a regular polygons interior angles is 156 degrees, how many sides does the polygon have?
------------
Ext angles = 180 - 156 = 24 degs
360/24 = 15 sides